### v2

@@ 5.1 Algorithm BL
Input: Preference matrix $\mathbf{A}$. Capacity vector m. Booking limit $\mathbf{b}$. $\gamma=\log_{1-\alpha}\delta$.
@@
Given any $\alpha,\delta\in(0,1)$,resource capacity $\mathbf{m}$ where $m_{j}=\Theta(m)$ for $j\in[L]$, and any input BL algorithm that attains a competitive ratio of $C$, GP-Enhanced-BL achieves the following: (i) is $(\alpha,\delta)$-fair, (ii) reaches a competitive ratio of $C-O\left(\frac{\log 1/\delta}{m}\right)$.
@@
$\inf_{I\in\mathcal{I}}\frac{\text{Rev}(\pi(I))}{\text{OPT}(I)}\geq\inf_{I\in\mathcal{I}}\frac{\text{Rev}(\textsf{BL}(I))-(\bar{a}+1)n\gamma\bar{r}}{\text{OPT}(I)}=C-O\left(\frac{\log 1/\delta}{m}\right).$
@@ 5.2 Algorithm NESTING
Initialize: Number of accepted customers $s_{i}=0$, $i\in[n]$. $\gamma=\log_{1-\alpha}\delta$.
@@
Given any $\alpha,\delta\in(0,1)$, and any input NESTING algorithm that attains a competitive ratio of $C$, GP-Enhanced-NESTING: (i) is $(\alpha,\delta)$-fair, (ii) reaches a competitive ratio of $C-O\left(\frac{\log 1/\delta}{m}\right)$.
@@ 6 Numerical Studies
In the second scenario, we consider a situation where individual fairness directly affects revenue. According to our survey detailed in Section 7, 56% of participants reported experiencing perceived individual unfairness in real. Thus, we model a scenario where a customer who is rejected, while a same type customer immediately before or after them is accepted, perceives this as unfair with a 56% likelihood. Reflecting the consequences of perceived unfairness, we integrate a revenue penalty based on our survey findings: 62% of such customers would stop purchasing from the store, and 69% would not recommend the store to others, signifying potential revenue loss. In this model, we reduce the total revenue by an amount $cr_{i}$ for each type $i$ customer who perceives unfairness, where $c$ represents the projected impact of the loss, essentially the expected frequency of visits by a type $i$ customer. For the purposes of our simulations, we set $c$ to values of $\{0.5,1,2\}$, to cover a range of possible loss intensities.
Results In the first scenario, we disregard individual fairness to assess its impact on revenue outcomes. As depicted in Figure 2, the discrepancy in total revenue between the two algorithms is marginal, adhering to the $O(\sqrt{T})$ boundary established in Theorem 4.1.
Figure 2:
@@
Figure 3:
@@ 7 Survey
The goal of our survey was to assess individuals’ buying behavior when they observed other customers receiving discounts of roughly 10% but they themselves did not. We recognize the significance of the specific discount value in this context; people’s valuation of money varies, and a discount perceived as negligible by some might not induce feelings of unfairness to them. To address this, we conducted a comparative survey, wherein the context incorporated a more substantial discount (specifically, 30%). This supplementary survey, which also comprised of 150 participants, is documented in Appendix 9.3. In all the scenarios we considered, we posit that there are two stores TrendThread and Stitch&Style which the survey respondent is considering. The products sold by the two places are identical in terms of quality and overall characteristics. The only difference is with regards to the pricing and the discounts that each stores offers.
@@ 8 Extension: Price-based Revenue Management
First, we employ the concept of a decreasing grace period, adjusting it to suit the price-based context. Specifically, a decreasing grace period to type $i$ customer within the interval $[t_{1},t_{2}]$ means that for every $u^{\text{th}}$ type $i$ customer arriving within the interval $[t_{1},t_{2}]$, we assign them a price $p_{u-1}^{(i)}$ with a probability of $1-\alpha$, and a price of +$\infty$+ ∞ with a probability of $\alpha$. This is analogous to the decreasing grace period for quantity-based RM.
@@
For the price-based revenue management problem where the price $p^{(i)}$ is given to type $i$ customers, by assigning a decreasing grace period $[t_{1}(i),t_{2}(i)]$, where $t_{1}(i)=\inf\{t:\min_{j\in[L]}m_{j}(t)\leq\bar{a}n\gamma\}$ and $t_{2}(i)=T$ to each type $i$ customer, where $\gamma=\log_{1-\alpha}\delta$, the $(\alpha,\delta)$-fair metrics is satisfied. Moreover, compared to the original static pricing algorithm, the revenue loss is bounded by $\frac{\bar{a}}{\underline{a}}n\gamma\bar{p}$, where $\bar{p}=\sup_{i\in[n]}p^{(i)}$.

### v3

@@ 6.1 Algorithm BL
Input: Preference matrix $\mathbf{A}$. Capacity vector m. Booking limit $\mathbf{b}$. $\gamma=\log_{1-\beta}\delta$.
@@
Given any $\alpha,\delta\in(0,1)$,resource capacity $\mathbf{m}$ where $m_{j}=\Theta(m)$ for $j\in[L]$, and any input BL algorithm that attains a competitive ratio of $C$, GP-Enhanced-BL achieves the following: (i) is $(\alpha,\delta)$-fair, (ii) reaches a competitive ratio of $C-O\left(\frac{\log_{1/(1-\beta)}1/\delta}{m}\right)$.
@@
$\inf_{I\in\mathcal{I}}\frac{\text{Rev}(\pi(I))}{\text{Rev}^{\star}(I)}\geq\inf_{I\in\mathcal{I}}\frac{\text{Rev}(\textsf{BL}(I))-(\bar{a}+1)n\gamma\bar{r}}{\text{Rev}^{\star}(I)}=C-O\left(\frac{\log_{1/(1-\beta)}1/\delta}{m}\right).$
@@ 6.2 Algorithm NESTING
Initialize: Number of accepted customers $s_{i}=0$, $i\in[n]$. $\gamma=\log_{1-\beta}\delta$.
@@
Given any $\alpha,\delta\in(0,1)$, and any input NESTING algorithm that attains a competitive ratio of $C$, GP-Enhanced-NESTING: (i) is $(\alpha,\delta)$-fair, (ii) reaches a competitive ratio of $C-O\left(\frac{\log_{1/(1-\beta)}1/\delta}{m}\right)$.
@@ 7 Numerical Studies
To implement GP-Enhanced-DLP, we evaluate the following parameter settings: $(\alpha,\delta)=\{(0.05,1/T),(0.05,1/\sqrt{T}),(0.3,1/T),(0.3,1/\sqrt{T})\}$. A smaller $\alpha$ corresponds to a more balanced allocation, enhancing fairness. Since (based on our definition of fairness) a fair algorithm ensures fairness with at least probability $1-\delta$, a smaller $\delta$ increases confidence in the fairness guarantee.
In the second scenario, we consider a situation where individual fairness directly affects revenue. According to our survey detailed in Section 8, 56% of participants reported experiencing perceived individual unfairness in real. Thus, we model a scenario where a customer who is rejected, while a same type customer immediately before or after them is accepted, perceives this as unfair with a 56% likelihood. Reflecting the consequences of perceived unfairness, we integrate a revenue penalty based on our survey findings: 62% of such customers would stop purchasing from the store, and 69% would not recommend the store to others, signifying potential revenue loss. In this model, we reduce the total revenue by an amount $cr_{i}$ for each type $i$ customer who perceives unfairness, where $c$ represents the projected impact of the loss, essentially the expected frequency of visits by a type $i$ customer. For the purposes of our simulations, we set $c$ to values of $\{0.5,1,2\}$, to cover a range of possible loss intensities. We also fix $(\alpha,\delta)=(0.05,1/T)$.
Results In the first scenario, we disregard individual fairness to evaluate its impact on revenue outcomes. As shown in Figure 2, the difference in total revenue between the two algorithms across all parameter choices $(\alpha,\delta)$ remains marginal, aligning with the $O(\sqrt{T})$ bound from Theorem 5.1.
Furthermore, the near overlap of curves corresponding to different parameter settings suggests that GP-Enhanced-DLP is robust to parameter choice. The intuition behind this low sensitivity is that in GP-Enhanced-DLP, $(\alpha,\delta)$ only influences the length of the grace period, given by $\log_{1-\alpha}\delta$. Even if this grace period is long—leading to a higher rejection rate of customers within a time segment $[(k-1)\sqrt{T},k\sqrt{T}]$—the algorithm compensates for this loss in the subsequent segment $[k\sqrt{T},(k+1)\sqrt{T}]$ by accepting an equivalent number of additional customers. As a result, any shortfall in one segment is recovered in the next. The only potential revenue loss occurs in the final segment or in the period when resources are fully depleted. However, this loss in one segment is negligible in comparison to the total revenue, confirming that GP-Enhanced-DLP maintains strong performance regardless of parameter selection.
Figure 2: Comparison between DLP-PA and GP-Enhanced-DLP (left: total revenue; right: regret).
@@
Figure 3: Comparison between DLP-PA and GP-Enhanced-DLP (left: $c=0.5$; middle: $c=1$; right: $c=2$).
@@ 8 Survey
The goal of our survey was to assess individuals’ buying behavior when they observed other customers receiving discounts of roughly 10% but they themselves did not. We recognize the significance of the specific discount value in this context; people’s valuation of money varies, and a discount perceived as negligible by some might not induce feelings of unfairness to them. To address this, we conducted a comparative survey, wherein the context incorporated a more substantial discount (specifically, 30%). This supplementary survey, which also comprised of 150 participants, is documented in Appendix 11.3. In all the scenarios we considered, we posit that there are two stores TrendThread and Stitch&Style which the survey respondent is considering. The products sold by the two places are identical in terms of quality and overall characteristics. The only difference is with regards to the pricing and the discounts that each stores offers. Additionally, at the survey’s conclusion, we inquired if participants had encountered or noticed such unfairness in the real world, and 56% of the respondents affirmed they had.
@@ 9 Extension: Price-based Revenue Management
First, we employ the concept of a decreasing grace period, adjusting it to suit the price-based context. Specifically, a decreasing grace period to type $i$ customer within the interval $[t_{1},t_{2}]$ means that for every $u^{\text{th}}$ type $i$ customer arriving within the interval $[t_{1},t_{2}]$, we assign them a price $p_{u-1}^{(i)}$ with a probability of $1-\beta$, and a price of +$\infty$+ ∞ with a probability of $\beta$. This is analogous to the decreasing grace period for quantity-based RM.
@@
For the price-based revenue management problem where the price $p^{(i)}$ is given to type $i$ customers, by assigning a decreasing grace period $[t_{1}(i),t_{2}(i)]$, where $t_{1}(i)=\inf\{t:\min_{j\in[L]}m_{j}(t)\leq\bar{a}n\gamma\}$ and $t_{2}(i)=T$ to each type $i$ customer, where $\gamma=\log_{1-\beta}\delta$, the $(\alpha,\delta)$-fair metrics is satisfied. Moreover, compared to the original static pricing algorithm, the revenue loss is bounded by $\frac{\bar{a}}{\underline{a}}n\gamma\bar{p}$, where $\bar{p}=\sup_{i\in[n]}p^{(i)}$.
10 Discussion
In this paper, we addressed fairness from the customer’s side and we provided a way for standard RM algorithms to satisfy our notion with almost no revenue loss. Defining the “correct” notion of fairness is a tall task, one which cannot be definitively solved by a single paper. That said, in an effort to move the discussion on fairness definitions in RM forward, we provide some more details around the properties of our fairness definition —and how we came to it— in this section.
10.1 Fairness Within and Across Customer Types
In Section 2, we established that in classical revenue management problems, customers are categorized based on the type and quantity of resources they request, rather than demographic characteristics or other personal attributes. In this setting, customers within the same type are homogeneous—they request the same resource(s) in the same amount and generate the same revenue—and our definition of individual fairness naturally applies within each type.
In real-world settings, however, it is common for different customer subgroups to request the same resource but generate different revenues. For example, airlines often sell the same economy-class seat at varying prices depending on the customer’s profile, and hotels may charge corporate clients lower rates than tourists for identical rooms. To model such cases, we extend the classical setting by allowing each type $i\in[n]$ to consist of $K$ subgroups, where each subgroup $k\in[K]$ requests the same resource in identical quantity but yields a different reward $r_{ik}$.
To maintain a unified framework, we show that this extended model can be equivalently transformed back into the classical network revenue management form. Specifically, let $\lambda_{ik}$ denote the arrival rate of subgroup $k$ within type $i$, satisfying $\sum_{k\in[K]}\lambda_{ik}=\lambda_{i}$. We define the weighted reward for type $i$ as
$\hat{r_{i}}=\frac{\sum_{k\in[K]}\lambda_{ik}r_{ik}}{\lambda_{i}}.$
By aggregating all subgroups into a single type using this weighted reward, we can treat them uniformly for the purposes of revenue optimization and fairness. As a result, the fairness mechanisms and algorithms developed in the main body of the paper can be directly applied by replacing the rewards of all subgroups within type $i$ with the aggregated weighted reward $\hat{r_{i}}$.
10.2 Time-Dependent Fairness: Ensuring Stronger Guarantees for Close Arrivals
Through our survey in Section 8, we examined how humans perceive fairness over time. Our findings indicate that people view disparate treatment occurring in close temporal proximity as significantly more unfair than similar disparities occurring in the distant past. This is intuitive: for instance, a customer is unlikely to care whether another customer receives preferential treatment 100 or 101 days later, whereas they would perceive a fairness concern if the disparity occurred today versus tomorrow.
To reflect this intuition, our fairness definition is time-dependent. Specifically, fairness should be locally strong—ensuring a low probability of observing disparate treatment for customers who arrive close in time—while allowing for weaker guarantees as the time gap increases. The three properties of $d(u,v)$ from Definition 3.1 make sure that we satisfy these desiderata. Property 1, which lets $d(u,v)$ be an increasing function with respect to $|u-v|$, ensures local fairness by imposing stricter constraints on customers who arrive closer together than those farther apart. Property 2 sets $d(u,v)=1$ when $|u-v|=1$, providing a consistent benchmark for evaluating fairness at the smallest time separation. Property 3 introduces a threshold $w_{\alpha}$ such that when $|u-v|=w_{\alpha}$, the fairness constraint becomes $\mathbb{P}\left[i(v)\succ_{\mathcal{A}}i(u)\right]\leq\alpha\cdot d(u,v)=\alpha\cdot 1/\alpha=1$, effectively placing no restriction on the algorithm. This reflects the idea that customers may not expect or require fairness comparisons with others who arrive significantly earlier or later. In practice, this property is reasonable—due to factors such as limited visibility into past transactions, changes in context or inventory, or simply diminishing concern over time, customers often do not compare their treatment to that of others far removed in time.
10.3 Incorporating Arrival Uncertainty: Fairness with High Probability
Our fairness definition is probabilistic. This is because directly enforcing fairness constraints in revenue management poses challenges due to the uncertainty in the arrival process. In particular, we will demonstrate next how such uncertainty can lead to highly inefficient allocation when fairness is imposed rigidly (i.e., with probability 1) by Equation (3). To illustrate this, we present a simple impossibility result: no algorithm that satisfies our fairness definition with probability 1 (i.e., $\delta=0$) can achieve a sub-linear regret even if there is only one type of customer!
Impossibility Result. Consider the simplest revenue management (RM) setting: a single type of customer arrives sequentially with rate $\lambda$, and a single type of resource is available with a total capacity of $M$ units. Each customer requests 1 unit of the resource and yields a revenue of 1 if accepted. Without fairness constraints, the optimal policy is trivial: accept every customer until the resource is fully depleted. There is no tradeoff—maximizing revenue simply means admitting as many customers as possible.
However, under the fairness constraint of Equation (3), the situation changes dramatically. No algorithm satisfying the fairness constraint can achieve a sub-linear regret (see Appendix 18). To see this, consider the case where only one unit of resource remains at some time $t$. If the algorithm accepts the arriving customer at time $t$, the next customer must be rejected due to depleted capacity. This immediate rejection creates a sharp disparity between two neighboring customers, violating the fairness constraint. As a result, any feasible algorithm must avoid fully depleting resources and must behave conservatively even when capacity is available.
To see the impact of this constraint on regret, consider the first customer whose acceptance probability is strictly positive. To satisfy the fairness constraint in Equation (3), the next customer must be accepted with probability at least $(1-\alpha)$ times that of the previous customer. This cascading requirement implies that acceptance probabilities cannot be reduced quickly. In fact, it takes infinite time steps to decay the acceptance probability from nearly 1 to 0, which means that the resource depletion must be earlier than the acceptance probability becoming 00. Hence, any algorithm satisfying the fairness constraint must set the acceptance probability of the first customer to be 00, which implies that the total revenue is 00.
This observation highlights a fundamental issue: enforcing fairness deterministically for every possible arrival sequence leads to highly inefficient allocation policies. In effect, the algorithm must behave as though the arrival process is adversarial—even when it is in fact stochastic—simply to avoid violating fairness in rare, extreme cases.
To illustrate this, consider again the basic revenue management setting with a single type of customer and one type of resource. Suppose the arrival rate is $\lambda=\frac{M}{2T}$, meaning that the expected number of customers over the horizon is only half the resource capacity. In most realizations of this process, the total number of arrivals will be well below $M$, and thus the algorithm could safely accept every customer while satisfying fairness. However, there is still a non-zero probability that more than $M$ customers arrive. Under the original fairness definition—which requires the constraint to hold for every realization—the algorithm cannot afford to accept all early customers, as doing so might unfairly exclude future ones in this rare event. So it must preemptively reject some customers even when the resource limit is not reached, simply to hedge against low-probability events.
To address this, we introduce an additional parameter $\delta$, which allows the fairness constraint to be enforced with high probability, rather than absolute certainty. Specifically, we require that Equation (3) holds with probability at least $1-\delta$. This formulation preserves fairness in the overwhelming majority of cases, while avoiding pathological behavior in rare scenarios. The parameter $\delta$ thus provides a principled way to balance fairness with efficiency in stochastic environments.